For a passenger who arrives at a certain bus stop at a random moment in​ time, the time spent waiting for the bus is uniformly distributed from 0 to 9 minutes. What is the probability someone who arrives at this bus stop at a random moment will wait at least 7 minutes for the​ bus? (Round to the nearest tenth of a​ percent.)
Answer #1

Solution :

Given that,

a = 0

b = 9

P(x  \geq 7 )

= (9 - 7) / (9 - 0)

= 2 / 9

=0.22

Probability = 0.22 %